John Vatalaro – Does the Vaccine Matter?

You have done a good job of explaining the issues raised by the Brownlee-Lenzer article and how they relate to things we have discussed in our class. There are a few points where you could say more (see comments below). Your writing is also quite good, although I think the production of the paper was a bit rushed and the proof-reading suffered as a result. I have marked some of the improvements there on the paper. Overall, though, this is good work.

Specific Comments

1. Technical point: The quotation reports a statement by Lisa Jackson, but you took it from the Brownlee-Lenzer article. In a case like this, the standard way to make the citation would be to identify who said the quotation in the text (that is, put Lisa Jackson’s name there). Then the citation would indicate where in the article the quotation was taken from and the parenthetical note (or the foot- or end-note) would identify the authors of the article.

2. I’m not an expert on the immunology here, but I think the concerns about mutation of the H1N1 virus are not directly related to the question of how many people build up immunity. I think there is some basic underlying mutation rate in the reproduction cycle of the virus that produces new genetic strains of the virus independently of whether people have immunity. A more important concern would be that the virus might mutate to acquire resistance to anti-viral medications like Tamiflu, etc.

3. As you say after this point, the situation is even more complicated than this because seasonal flu is always caused not just by one virus strain, but by several. (Also, in cases like this where there is a common word like “strain” that is being used in a technical sense, it is not a good idea to try to find synonyms to vary the word choice. “Strand” really means something different. It’s OK to use “strain” every time even if it sounds like you are repeating yourself.)

4. “Adds onto” does not really sound right here. Maybe: “Influenza also interacts with sicknesses such as AIDS, asthma, and cancer to cause additional deaths.”

5. This is one place where what you say is not actually correct. It’s not that Brownlee and Lenzer are “arguing” that the cohort studies have these fantastic results. They are reporting the claims made by those studies to call them into question. I know you know this, but what you wrote does not make this clear.

6. As the article reports, there is pretty good evidence that the seasonal flu vaccine produces a good immune response in young, healthy people. So in fact, this seems to me to be one of the strongest arguments for giving the vaccines. If you can prevent the spread of the virus to the people who are most at risk, then that is a good thing.

7. You say medical professionals are “not too fond” of placebo-controlled trials for flu vaccines “for ethical reasons.” That is accurate. But you seem to be overlooking the main reason for the strength of their objections. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, aren’t you doing harm if they get sick and possibly die from the illness? This seems to be a real ethical dilemma to me!

8. This is the wrong word – you mean “efficacy” or ”effectiveness” (that is: whether vaccines work at all, not how well they work).

Content:

Form:

Brandon Nunn – “Does the Vaccine Matter”

You have done a very good job of explaining the issues raised by the Brownlee-Lenzer article and how they relate to things we have discussed in our class. There are a few points where I think there are additional issues that you have not accounted for, or are dismissing too quickly (see comments below). But over all, this is strong work. Your writing is also generally very good, although there are a few questionable word choices.

Specific Comments

1. The “healthy user” effect could also apply when healthy elderly people are vaccinated, but sicker (say home-bound or bed-ridden) elderly people are not able to get to vaccination centers. That is, it’s not just the age of the vaccine recipients (or non-recipients) that matters. It’s really their general condition.

2. “Connoisseur” is not the right word here. I think you mean “expert.” One can be a “connoisseur” of opera, fine wines, or gourmet food, but not of vaccines(!) Beware of thesaurus suggestions for synonyms. There are very few cases where two words are exact equivalents. Words generally have shades of meaning that make all the difference.

3. You say medical professionals have ethical concerns about placebo-controlled trials for flu vaccines. That is accurate, although I don’t think that the main reason is concern for their reputations. You seem to be overlooking the main reason for the strength of their objections. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, aren’t you doing harm if they get sick and possibly die from the illness? To justify that action, you have to argue that the greater good for society of doing the vaccine trial outweighs the concerns of individual patients. But that is not the way doctors are trained to think and I am not sure I would want their focus on the individual to change(!) This seems to be a real ethical dilemma with no easy answers to me.

4. This is a very good point. The “flip side” is that stories about courageous pioneers in science “bucking the received wisdom” and staking out controversial new positions have a certain appeal for some people. Some of us like to “root for” the underdogs in this kind of situation. I think that angle of the story is probably one reason why this article appeared in a general-audience periodical like the Atlantic.

Content:

Form:

Kyle Walraven – No title

You have done a good job of explaining the issues raised by the Brownlee-Lenzer article and how they relate to things we have discussed in our class. There are a few points where you could say more (see comments below). Your writing is also good (but see the suggestions for changes written into the paper). Overall, this is very good work.

Specific Comments

1. You say medical professionals have ethical concerns about placebo-controlled trials for flu vaccines. That is accurate, but you almost seem to be minimizing the main reason for the strength of their objections. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, aren’t you doing harm if they get sick and possibly die from the illness? To justify that action, you have to argue that the greater good for society of doing the vaccine trial outweighs the concerns of individual patients. But that is not the way doctors are trained to think and I am not sure I would want their focus on the individual to change(!) This seems to be a real ethical dilemma with no easy answers to me.

2. This is a good point. There is undoubtedly a certain appeal connected to stories about courageous pioneers in science “bucking the received wisdom” and staking out controversial new positions. I think that angle of the story is probably one reason why this article appeared in a general-audience periodical like the Atlantic.

3. The discussion of what a non-expert can do in the circumstances is good, although I think you are being somewhat too pessimistic. If you believe the model of science that Grinnell proposes, the credibility process will tend to produce more and more reliable results over time. So the situation is not always going to be what it is now where even the “experts” disagree.

Content:

Form:

Mary Arnold – “Does the Vaccine Matter?”

You have done a very good job of explaining the issues raised by the Brownlee-Lenzer article and how they relate to things we have discussed in our class. There are a few points where you could say more (see comments below). Your writing is also very good (I only have a few suggestions for changes written into the paper). Overall, this is very good work.

Specific Comments

1. If you go back and look again at the Brownlee-Lenzer article, you will see that Jefferson was trained at the London School of Tropical Hygiene. The article says he is “currently a physician based in Rome and the head of the Vaccines Field at the Cochrane Collaboration.”

2. As you point out later, this is not the whole story, though. One of the reasons to immunize even the young and healthy is so that they do not spread the virus to older and sicker individuals whose immune systems are less able to protect them.

3. This is the point I was making in 2. above, and it just might be one of the strongest arguments in favor of the flu vaccinations. Studies have shown pretty conclusively that the vaccines do produce strong immune responses (production of antibodies) in younger, healthier people. If they are not as susceptible to infection, they may also not be spreading the virus around to people who are more susceptible to the virus and secondary infections.

4. This is an interesting secondary aspect of this article. There is undoubtedly a certain appeal connected to stories about courageous pioneers in science “bucking the received wisdom” and staking out controversial new positions. I think that angle of the story is probably one reason why this article appeared in a general-audience periodical like the Atlantic.

5. A final general comment. You say medical professionals have ethical concerns about placebo-controlled trials for flu vaccines. That is accurate, it would be good to say some more about the main reason for the strength of their objections. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, aren’t you doing harm if they get sick and possibly die from the illness? To justify that action, you have to argue that the greater good for society of doing the vaccine trial outweighs the concerns of individual patients. But that is not the way doctors are trained to think and I am not sure I would want their focus on the individual to change(!) This seems to be a real ethical dilemma with no easy answers to me.

Content:

Form:

Andrew Coury – “The H1N1 Vaccine: a False Sense of Security”

You have done a good job of explaining the issues raised by the Brownlee-Lenzer article and how they relate to things we have discussed in our class. There are a few points where you could say more or where what you say seems to be a bit questionable (see comments below). Your writing is also good (but see the suggestions for changes written into the paper). Overall, this is good work.

Specific Comments

1. The end of the previous paragraph and the start of the next one at this point appear to be contradicting one another. I think the first statement reflects initial estimates of the possible severity of an H1N1 outbreak. If those estimates came true, then the 90,000 deaths would not be “a relatively small number.” It would be significantly larger than (2.5 times, in fact) the usual 36,000 deaths per year attributed to seasonal flu. It is true that other causes of death produce greater mortality. But this would still be a rather large number of deaths compared to the expected levels.

2. You are referring to what is often called the “healthy user effect.”

3. In fact the kind of randomized, placebo-controlled trials that Jefferson is proposing have never “gotten off the ground.” So saying he “introduce[d]” one is not correct.

4. You say medical professionals have resisted placebo-controlled trials for flu vaccines. That is accurate, it would be good to say some more about the main reason for the strength of their objections. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, aren’t you doing harm if they get sick and possibly die from the illness? To justify that action, you have to argue that the greater good for society of doing the vaccine trial outweighs the concerns of individual patients. But that is not the way doctors are trained to think and I am not sure I would want their focus on the individual to change(!) This seems to be a real ethical dilemma with no easy answers to me.

5. “Chaos would erupt” seems somewhat overstated to me. Of course, I cannot really predict what the reaction would be in that situation, but I think this kind of breakdown in society would only happen if the death rate from the H1N1 flu was extremely high, and the vaccines were shown to be ineffective, and maybe there was some flagrantly improper behavior on the part of the vaccine producers and the government that came to light. Whether or not the vaccines work the way most doctors think they do, I don’t know any credible evidence that the push to produce the vaccines and the antivirals has been improper (that is, ethically shady). Raising questions is not the same as proving misdeeds!

6. You’re getting close to the point I was making in item 4 above. But if you look at it from the point of view of a practicing physician, the issues become, if anything, even more confusing(!)

Content:

Form:

Paige Carr – What Should We Do? Questioning the Flu Vaccine

Your discussion of the appeal that “maverick” researchers like Lisa Jackson and Tom Jefferson have for readers is especially good (see comments 3 and 4 below). The rest of the paper is also a good summary of some of the issues raised by the Brownlee-Lenzer article and how they related to things we have discussed in this course. The writing seems a bit technically rough this time, though (see the comments and corrections on the paper itself). There are a few instances where you used the wrong word. There are also some awkward or ungrammatical sentences. I think this might have been rushed because of the end of the semester. But be sure you put in some more time on proof-reading in the future.

Specific Comments

1. This sounds very cynical. Are you saying that one cannot believe anything one hears in the news? While there are certainly efforts to falsify the record and spread incorrect information, is everything suspect?

2. The views of the Lisa Jackson, Tom Jefferson, and other scientists reported in this article are very much minority views at present. So it is important to qualify this by saying “some flu researchers.”

3. I think you have “nailed” the “good story” aspect of this article. There is undoubtedly a certain appeal connected to stories about courageous pioneers in science “bucking the received wisdom” and staking out controversial new positions. I think that angle of the story is probably one reason why this article appeared in a general-audience periodical like the Atlantic.

4. See the previous comment. Portraying these scientists as such underdogs in the “credibility process” makes them sympathetic figures for readers.

5. You say medical professionals have ethical concerns about placebo-controlled trials for flu vaccines. That is accurate, it would be good to say some more about the main reason for the strength of their objections. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, aren’t you doing harm if they get sick and possibly die from the illness? To justify that action, you have to argue that the greater good for society of doing the vaccine trial outweighs the concerns of individual patients. But that is not the way doctors are trained to think and I am not sure I would want their focus on the individual to change(!) This seems to be a real ethical dilemma with no easy answers to me.

6. Let me play “devil’s advocate” for a moment and bring up the following question. Suppose you decide after consulting your conscience that you will not get the flu vaccine. But suppose you are a flu “carrier” who does not get sick but spreads the virus around to others who have weaker immune systems and get sick from the virus. Is that something that can be justified? Doesn’t society have an interest in protecting its more vulnerable members from this situation? In fact, to me one of the strongest arguments for vaccination, even with the all the questions, is that it may do something to prevent the transmission of the virus by boosting the immune responses of the young, healthy members of the population.

Content:

Form:

Jonathan Ye – Is Over-Preparing for the Flu Possible?

This is a mostly good discussion of some of the issues about influenza vaccines raised by the Brownlee-Lenzer article. Your way of looking at the situation is probably a prudent one, especially if a really virulent strain of influenza emerges in the future (or H1N1 mutates into something much worse than it now appears to be). I think one omission, though, is that you only tangentially address what is probably the main thrust of what they are saying – namely, that double-blind, randomized, placebo-controlled trials of flu vaccines are needed to determine whether they really work as advertised. See item 2 below for the main reason why this has not happened and some of the ethical questions involved.

Specific Comments

1. This makes it sounds as though the media were making up alarming figures about swine flu deaths. I think they have tended to sensationalize in some cases. But also recall that the early projections of how serious the swine flu pandemic would be (from the CDC and other scientists) were much worse-sounding than the way things have actually turned out (so far, at least!) Some of the original projections had something like 90,000 deaths in the US. It does not look anything like that serious now, of course, so it is easy to fault some of the early reporting. But at the time, I am sure that the media just thought they were doing their part to educate people to a threat to public health.

2. You say correctly that double-blind randomized trials with use of placebos would probably yield the most conclusive evidence about whether flu vaccines are effective. But as Brownlee and Lenzer indicate no such trials have ever been carried out, and they are not likely to be carried out. You don’t mention a key point about why this is the case: physicians have major ethical concerns about placebo-controlled trials for flu vaccines. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, aren’t you doing harm if they get sick and possibly die from the illness? To justify that action, you have to argue that the greater good for society of doing the vaccine trial outweighs the concerns of individual patients. But that is not the way doctors are trained to think and I am not sure I would want their focus on the individual to change(!) This seems to be a real ethical dilemma with no easy answers to me.

3. I am not sure that I “buy” your reasoning for why you think the estimate that as many as 50% of health-care professionals are reluctant to be vaccinated for flu is incorrect. It might not be anywhere near that high, but I think your objections are mostly wide of the mark. It is certainly true that not all health-care professionals have direct contact with patients. But the possibility of passing or spreading the virus due to contacts with patients is not the only reason why such professionals might consider getting the vaccine. If they believed in the effectiveness of the vaccine, wouldn’t getting the vaccine be a good idea simply to protect their own health? Also, the issue is not whether these professionals believe in vaccines in general. I’m sure almost all of them do because of the success of vaccines in eradicating other diseases like smallpox, polio, etc. The issue is really whether the seasonal and H1N1 flu vaccines do what their supporters think they do. And there, I think that many health-care professionals think the jury is still out (even if they are not as negative about the vaccines as some of the people mentioned in the article are). Finally, although it has not received much media coverage, I know that there have been some cases where health-care professionals have resisted mandates that they get the vaccines. (This has probably remained “under the radar” because the availability of H1N1 vaccine has been so limited so far, and the available doses have been reserved for individuals considered to be at high personal risk – pregnant mothers, younger children, etc.)

Content:

Form:

Stephanie Craig – Immunization: Is It Really All It Is Cracked Up To Be?

At the end, you come down pretty squarely in favor of randomized, double-blind, placebo-controlled clinical trials for influenza vaccine. I’m wondering how you would answer the questions raised in point 2 below, though. Do you believe that the benefit to society as a whole of knowing better evidence of the efficacy (or not) of flu vaccine outweighs the potential harm that might be caused to participants in the trials? Would getting some real form of “informed consent” from the participants possibly address this objection?

In any case, your paper is well-written and makes some good connections with topics we discussed this semester. Good work.

Specific Comments

1. You are right to point out that this view is based on the underlying assumption that the vaccine does reduce the chance of death or other serious complications from the flu. But it would be good to say some more about the main reason for the strength of doctors’ objections to doing this kind of trial. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, aren’t you doing harm if they get sick and possibly die from the illness? To justify that action, you have to argue that the greater good for society of doing the vaccine trial outweighs the concerns of individual patients. But that is not the way doctors are trained to think and I am not sure I would want their focus on the individual to change(!) This seems to be a real ethical dilemma with no easy answers to me.

2. “Widespread panic would ensue” seems somewhat overstated to me. Of course, I cannot really predict what the reaction would be in that situation, but I think this kind of breakdown in society would only happen if the death rate from the H1N1 flu was extremely high, and the vaccines were shown to be ineffective, and maybe there was some flagrantly improper behavior on the part of the vaccine producers and the government that came to light. Whether or not the vaccines work the way most doctors think they do, I don’t know any credible evidence that the push to produce the vaccines and the antivirals has been improper (that is, ethically shady). Raising questions is not the same as proving misdeeds!

Content:

Form:

Josh Hauser – Swine Flu

Your paper is OK as a summary of some of the issues raised in the Brownlee-Lenzer article. But it never goes into anything in much depth and there are some important points that are glossed over or misstated (see comments below). It seems to me that you could have put some more time and effort into this.

Specific Comments

1. This is an overstatement if you take the whole world into account. If you are talking about the US, and the rest of the industrialized world, then this is closer to being true. Be more careful about the broad, general statements like this.

2. The first flu vaccines were developed in the 1940’s and widespread use of influenza vaccine is actually a relatively recent development (the last 50 years or so). When I was a child, for instance, use of such vaccines was not promoted to anything like the degree that is true now. Also, the first sentence of the second paragraph just repeats something you said in the previous paragraph.

3. You say correctly that double-blind randomized trials with use of placebos would probably yield the most conclusive evidence about whether flu vaccines are effective. But as Brownlee and Lenzer indicate no such trials have ever been carried out, and they are not likely to be carried out any time soon. You don’t mention a key point about why this is the case: physicians have major ethical concerns about placebocontrolled trials for flu vaccines. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, aren’t you doing harm if they get sick and possibly die from the illness? To justify that action, you have to argue that the greater good for society of doing the vaccine trial outweighs the concerns of individual patients. But that is not the way doctors are trained to think and I am not sure I would want their focus on the individual to change(!) This seems to be a real ethical dilemma with no easy answers to me.

4. I am not an expert here, but I don’t think immunologists would say that that viruses can become “resistant to vaccines.” The vaccine just produces an immune response in the vaccinated person (production of antibodies). That result provides some protection if the person later encounters a similar virus. A virus could later mutate to become less susceptible to the person’s augmented immune system, but that is not resistance to the vaccine itself. On the other hand, the development resistance to antiviral drugs by viruses is a serious concern.

5. Careful – it is not Tamiflu itself that is the confounding factor. The confounding factor is that the people who can afford Tamiflu have higher standards of living, better living conditions, better hygiene, better health care, etc. So the confounding factor is actually the higher income level. (It is associated both with access to Tamiflu and to the outcome of treatment for flu.)

6. There is another side of the “good story” aspect of this article. There is undoubtedly a certain personal appeal in stories about courageous pioneers in science “bucking the received wisdom” and staking out controversial new positions. I think that angle of the story is probably one reason why this article appeared in a general-audience periodical like the Atlantic.

Content:

Form:

Patrick Jarvis – Does the Vaccine Matter?

Patrick, your paper about the Brownlee-Lenzer article is well written. Your discussion of the reason why no randomized, double-blind, placebo-controlled studies of flu vaccines have been done (page 2 of your paper) is also very good. However, based on things here, and on the first paper, I think we really need to work on your close reading skills. You seem to jump to unwarranted conclusions at times (perhaps because of your own personal opinions). Some of the things you are saying here are serious misrepresentations of what Brownlee and Lenzer are really getting at (see comments 2 and 4 below especially). Either because of how you read their article, or because of a lack of care in writing, you are off the mark in several places.

Specific Comments

1. Strictly speaking, of course, the virus did not shut down Mexico City or colleges. The closures were precautionary measures by public health officials to prevent even wider spread of the virus. This is an important point because it shows that many of the inconvenient effects of this pandemic have actually been the results of choices made by humans. On the other hand, I don’t want you to think I am arguing that those closures and other measures were unnecessary. If the H1N1 strain this year had turned out to be as bad as the 1918 - 1919 “Spanish flu,” the situation would have been very different. The problem is that there is no way to tell ahead of time whether a new flu strain is mostly a nuisance (like this year) or a major public health disaster (like 1918 - 1919).

2. I think this is a crucial misreading of their main point. (You are not alone in thinking this, but I am sure they would disagree pretty strenuously with what you are saying!) If you read more carefully, you will see that Brownlee and Lenzer never say they do not think people should get the vaccine. They say that it has never been satisfactorily demonstrated that the flu vaccine works the way its supporters claim. And, they have other questions about the way government health officials have “pushed” the vaccine and anti-viral drugs even though there is no real proof that they work. Especially if there was another major public health disaster like the 1918 - 1919 “Spanish flu,” their point is that we have to be better prepared (perhaps by producing better vaccines or by developing other measures).

3. See point 1 above. The problem is that there is simply no way to tell ahead of time how severe a public health challenge a new influenza strain will be. In fact, there may not even be any way to tell (even in principle), because there is an essential component of randomness involved in the genetic mutations that produce new strains of flu virus.

4. Careful again – the mocking is most definitely not directed against the vaccine itself. It is directed against the claim made in the “cohort studies” that immunization could reduce deaths by 50% to 90%. This is another important place where I don’t think your reading of what the article was saying was careful enough.

Content:

Form:

Nick Foster – (No title)

Your discussion of the ethical issues that are cited by doctors as the reasons for why no randomized, double-blind, placebo-controlled trials of flu vaccines have been undertaken is very good. You have also summarized most of the other points in the Brownlee-Lenzer article very well and raised some interesting questions. Your writing is generally very good, although there are a few unclear places and minor slips. (See the corrections and suggestions on the printout.)

Specific Comments

1. Using “find” and “discover” sounds strange here because it sounds as though the vaccines exist somewhere out there and we just have to locate them. It would be better to use “develop” or “produce” because what is actually being done is taking general methods for producing vaccines and adapting them to the particular strains of flu virus that are the targets at this time.

2. There seems to be a typo or something missing here. I cannot tell what this is supposed to be saying.

3. “Knowing for sure” is probably too strong here. There is always some question about whether a scientific result is completely credible. Those studies would make the evidence for the effectiveness of flu vaccine much stronger, but not certain.

4. There is also an interesting secondary aspect of this article. There is undoubtedly a certain appeal connected to stories about courageous pioneers in science “bucking the received wisdom” and staking out controversial new positions. If they turn out to be correct in the long run, they are often the individuals looked on as heroes in the subsequent history of the subject. I think that angle of the story is probably one reason why this article appeared in a general-audience periodical like the Atlantic.

Content:

Form:

Dan Schmader – Does the Vaccine Matter?

This is a reasonably good summary of the ideas from the Brownlee-Lenzer article. You might have brought out connections with things we discussed in class somewhat more explicitly in some places. The writing is also general good on a “local level.” But the paper as a whole is somewhat repetitive. I think making a tight outline before starting to write might have helped here.

A minor technical point: it’s a number of people, not an amount of people. People come in individual units. You use “amount” for things like flour or liquids where it is difficult or impossible to isolate individual units of “stuff.”

Specific Comments

1. This is what is commonly called the “healthy user bias.” It is a classic example of a possible confounder in studies of the effectiveness of a medical treatment.

2. This is again an analyis of possible confounders in the “cohort studies” commonly used to justify the effectiveness of the flu vaccines. It would be good to make the connection with the topics we discussed in class that are very closely related.

3. It is true that there is a certain amount of reluctance to accept new ideas on the part of doctors. But there is another side of the “good story” aspect of this article too. There is undoubtedly a certain personal appeal in stories about courageous pioneers in science “bucking the received wisdom” and staking out controversial new positions. I think that angle of the story is probably one reason why this article appeared in a general-audience periodical like the Atlantic. the “good story” aspect of this article.

4. Grinnell’s book is definitely not a novel. A novel is a work of fiction in a long format. That does not apply to “Everyday Practice of Science.”

5. You say medical professionals have ethical concerns about placebo-controlled trials for flu vaccines because of the possibility of someone dying in the trial. If you dig into this a little deeper, though, things get even more difficult to decide. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, aren’t you doing harm if they get sick and possibly die from the illness? To justify that action, you have to argue that the greater good for society of doing the vaccine trial outweighs the concerns of individual patients. But that is not the way doctors are trained to think and I am not sure I would want their focus on the individual to change(!) This seems to be a real ethical dilemma with no easy answers to me.

Content:

Form:

Greg Monfette – The Pandemic-Preventing Vaccine’s Legitimacy

This is a reasonably good summary of the points raised in the Brownlee-Lenzer article. You have done a good job, too, of connecting their points to things we discussed in class. (However, I think the organization of the paper somewhat undercuts the connections – see point 5 below.) The writing is very good, with a few relatively minor exceptions (see the corrections and suggestions marked in the printout).

Specific Comments

1. The Atlantic Monthly is actually a very old print magazine (first published about 150 years ago). It has only recently “gone online.”

2. “Unraveled” is a strange word choice here. I think you mean devastated or swept through, or something like that. Also, it is important to realize that the 1918-1919 “Spanish flu” epidemic was a world-wide phenomenon, not limited to the US.

3. I think you are referring her to the “cohort studies” of the seasonal flu vaccines that are commonly used to justify the vaccines’ effectiveness. The claim is actually that the vaccines reduce mortality (deaths) by 50%. Critics like Tom Jefferson are mostly critical of this kind of claim which is pretty clearly too strong.

4. Saying that “many of the top researchers question the validity of this” is not accurate. People like Lisa Jackson, Tom Jefferson, etc. mentioned in this article are actually pretty rare and definitely outside the mainstream of the medical community’s opinions about the usefulness of the flu vaccines.

5. In terms of the organization of the paper, I think it would have been better to combine the discussion you have on page 4 with this point. Here you are just saying there are confounding factors, but there you are explaining what they are and how they relate to what we discussed. It would be more unified not to break things up the way you did.

6. You say medical professionals have ethical concerns about placebo-controlled trials for flu vaccines because of the possibility of someone dying in the trial. If you dig into this a little deeper, though, things get even more difficult to decide. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, aren’t you doing harm if they get sick and possibly die from the illness? To justify that action, you have to argue that the greater good for society of doing the vaccine trial outweighs the concerns of individual patients. But that is not the way doctors are trained to think and I am not sure I would want their focus on the individual’s health to change(!) This seems to be a real ethical dilemma with no easy answers to me.

7. “Worldwide fear and chaos” seems somewhat overstated to me. Of course, I cannot really predict what the reaction would be in that situation, but I think this kind of breakdown in society would only happen if the death rate from the H1N1 flu was extremely high, and the vaccines were shown to be ineffective, and maybe there was some flagrantly improper behavior on the part of the vaccine producers and the government that came to light. Whether or not the vaccines work the way most doctors think they do, I don’t know any credible evidence that the push to produce the vaccines and the antivirals has been improper (that is, ethically shady). (Raising questions is not the same as proving misdeeds!)

Content:

Form:

Carl Losito – Does the Vaccine Matter?

The technical aspects of your writing show some improvement here. However, I am afraid that what you are saying, while clearer, is often confused, incorrect, or incomplete. See the comments below for some major examples of this.

Specific Questions

1. This first paragraph is somewhat incoherent because it mixes the discussion of seasonal flu and H1N1 flu to such an extent. It also overdramatizes the seriousness of H1N1 flu according to the data that has come out over the past several months. If the current patterns hold, the death rate from H1N1 flu will actually be much lower than originally predicted, and even lower than a typical seasonal flu.

2. There may be some people who oppose all vaccines for the reasons you state here. However, that is certainly not the issue raised by Brownlee and Lenzer in this article. Their concern is that we are relying too much on flu vaccines that have never been convincingly proved to work as well as they are supposed to. If a really serious flu epidemic comes (like the 1918 - 1919 “Spanish flu”), they want to warn people that we might be putting too much confidence in these unproven vaccines.

3. At least so far, this is not really accurate. H1N1 infections have not been especially severe or more serious than the usual seasonal flu strains.

4. This was one of the very early predictions that was made before the outbreak really started in the US. Unless the virus mutates in a way to make it much more virulent, I doubt that anyone thinks that many deaths will actually result in the US.

5. There is a somewhat subtle point here that I think you are not appreciating. The seasonal flu typically kills more older people than younger ones (partly because there are many older, sicker individuals who succumb to secondary infections, etc.) Most of the efforts to immunize against seasonal flu have targeted this age group. But the situation is reversed for the H1N1 flu. Most of the more severe cases and deaths (for whatever reason) have occurred in younger patients, and young children, pregnant women, etc. are the highest priority recipients of the H1N1 vaccine as it comes on line. So what you are saying here is really only applicable to the H1N1 flu vaccination effort. The article is dealing with both H1N1 and seasonal flu, though, so it matters which vaccination effort you are talking about.

6. Careful – antibiotics are drugs that target bacteria. Influenza is a viral disease and treatments for it involve a combination of vaccination and anti-viral drugs, which are different.

7. You really need to qualify this statement. This point of view is actually held by only a small minority of experts (many of whom are quoted in this article). Most immunologists believe that the “cohort studies” questioned by people like Lisa Jackson, Tom Jefferson, and Sumit Majumdar have demonstrated the effectiveness of seasonal flu vaccines.

8. Careful – confounding factors can lead to misleading results in a study. They cannot affect a vaccine’s value – it either works to the desired degree or it does not.

9. The concluding sentences of this paragraph are very confused. The adverse reactions you mention are ones that have been reported with Tamiflu. But Tamiflu is not a vaccine. It is an anti-viral drug that is supposed to help an infected person combat the virus when it is already present in his or her body. Then, the moral and ethical arguments are usually made against the idea of doing double-blind, randomized, placebo-controlled studies of vaccines or drugs like Tamiflu. The kind of clinical trials you are describing have, in fact, never been done. You just throw out the phrase “moral and ethical reasons” without explaining what the issues are. This is an important point that deserves a lot more discussion. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, aren’t you doing harm if they get sick and possibly die from the illness? To justify that action, you have to argue that the greater good for society of doing the vaccine trial outweighs the concerns of individual patients. But that is not the way doctors are trained to think and I am not sure I would want their focus on the individual’s health to change(!) This seems to be a real ethical dilemma with no easy answers to me.

10. The issue is not whether technology will become available. It is whether the trials can be done in a way that does not raise the ethical issues raised in the previous point.

Content:

Form:

Amanda Cabrera – (No Title)

You were the only one in the class who chose this topic! I am glad you did, though, because your paper does a very good job of describing Strogatz’s blog post about looking at the world through mathematical eyes, and the specific pattern given by Zipf’s Law. Your writing is generally very good too!

Specific Comments

1. Don’t overuse the word “unique.” This should be applied only to cases where there really is only one of the things you are talking about. I think what you mean here is closer to “uncommon” or “unusual.”

2. It is good that you mentioned this. And of course, many of the critical comments about Strogatz’s blog post were pointing out exactly this aspect of the pattern (that the largest cities can be different). There is also a subtler point here, though. Namely, Zipf’s Law is, like most scientific patterns, subject to measurement errors and “noise” in the data. This means that the Population = c(rank)^{k} power law is never going to match the measurements of populations for actual cities exactly. The pattern only becomes apparent if you do the kind of statistical data analysis we did in the lab (and you look at enough of the cities to see whether the largest ones are outliers to the main pattern). Of course, that kind of detailed analysis would be difficult to describe in a general-audience periodical’s blog where authors cannot assume that readers have the mathematical background to appreciate the procedures carried out in the data analysis. In a way, Strogatz was forced to leave out most of the stuff that would have addressed the objections the readers came up with. (It is really hard to write correctly and meaningfully about math or science for a general audience!) But I agree that he might have “headed off” some of the criticism by saying somewhere that Zipf’s Law is really an approximate relation, not an equation that holds exactly in all possible cases, and the pattern holds better for cities farther down in the rank hierarchy. He was leaving himself open to silly responses by not making that aspect of the topic clearer.

3. You could view this as a quantitative measure of “economies of scale” (as discussed in economics)!

4. I think your objection here can be attributed the fact that you are looking at living organisms on an inappropriate distance and size scale (thinking of the whole rather than the parts, in effect). Think about what a living body looks like on the level of individual cells, blood vessels, etc. At that scale, doesn’t a living body start to resemble a city with smaller streets (capillaries), major highways (big arteries and veins), “traffic” (blood flow), utilities (organs), etc.(!)

5. Maybe, but see point 2 above. Telling the whole story requires readers to have some knowledge and tolerance for mathematics that is not always present!

Content:

Form:

Shelby Proft – (No Title)

Your writing is much improved this time (see the comments on the printout for some suggestions and corrections). You are entitled to your own opinion on this issue, and there are certainly a lot of people who would agree with the stand you are taking here. However, I don’t think the Brownlee-Lenzer article is really making the same point. It’s quite a bit more subtle (see comments 1 and 5 below especially!)

Specific Comments

1. I have to disagree pretty strongly with this statement. (And Brownlee and Lenzer, the authors of the article, would disagree too!) There have been a lot of the “cohort studies” discussed in the article done that appear to support the claim that seasonal flu vaccine reduces mortality from flu. The skeptics mentioned in the article (Lisa Jackson, Tom Jefferson, etc.) have publicized some contrary results, but their work is actually still very controversial in the medical community. Moreover, Brownlee and Lenzer’s main point, as I see it, is not that the flu vaccines are useless. Instead, they want to raise these issues because the methodological problems with the cohort studies that are the main evidence for the flu vaccine supporters mean that (in their view) we don’t really know whether flu vaccines work “as advertised.” They may work just fine, but we don’t know because the correct tests have not been done to decide! This is especially serious because most public health people would agree that it is only a matter of time before a really serious flu epidemic (like the 1918 - 1919 “Spanish flu,” not like H1N1 this year, apparently, at least so far) occurs again. Hence the concern they point out about a “false sense of security.”

2. Since you have not mentioned Lisa Jackson before or explained which side of the issue she has argued, it would be good to put in a sentence introducing all of that before this quotation.

3. As the article reports, there is pretty good evidence that the seasonal flu vaccine produces a good immune response in young, healthy people. So in fact, this seems to me to be one of the strongest arguments for giving the vaccines. If you can prevent the spread of the virus to the people who are most at risk by lowering the incidence of the disease in younger, healthier people who might spread it around, then that is a good thing.

4. Careful – these side effects are actually reported with use of Tamiflu, not the seasonal flu vaccine. The effects (and effectiveness) of vaccines and of anti-virals are two different issues!

5. The most perplexing aspect of this whole story is probably why medical professionals are opposed to placebo-controlled trials for flu vaccines “for ethical reasons.” You do not say enough about this, but it is a key point in this whole discussion. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, haven’t you done harm if they get sick and possibly die from the illness? To justify that action, you have to argue that the greater good for society of doing the vaccine trial outweighs the concerns of individual patients. But that is not the way doctors are trained to think and I am not sure I would want their focus on the individual to change(!) This seems to be a real ethical dilemma to me!

Content:

Form:

Jack Karlovec – Does the Vaccine Matter?

This is reasonably good as a summary of the start of the Brownlee-Lenzer article. But you spend too much time at the start describing the MedImmune factory procedures, which are of really limited relevance. The rest is somewhat superficial and you have only touched on some of the key points. Plus, I think you have misunderstood what might be the key point of the article. See comments 2 and 4 below. Your writing is generally better than on the first paper. But it is not considered good writing style to break out points in a “bullet list” (like the one on your page 3) in this kind of writing. Use complete sentences and paragraphs. The “bullet lists” are fine for oral presentations or more informal memos, etc. But don’t use them here.

Specific Comments

1. These figures are actually the averages for years in which only the “usual,” seasonal flu viruses are presesent. The original estimates were that H1N1 could kill as many as 90,000 people in the US this year. Fortunately, it now seems unlikely that the total will be that high. But there is really no way to predict how serious a new flu strain will be until after its effects are observed.

2. You seem to be making a crucial misreading of the main point of the Brownlee-Lenzer article. (You are not alone in thinking this, but I am sure they would disagree pretty strenuously with what you are saying!) If you read more carefully, you will see that Brownlee and Lenzer never say they do not think people should get the vaccine. They say that it has never been satisfactorily demonstrated that the flu vaccine works the way its supporters claim it does. And, they have other questions about the way government health officials have “pushed” the vaccine and anti-viral drugs even though there is no real proof that they work. Especially if there was another major public health disaster like the 1918 - 1919 “Spanish flu,” their point is that we have to be better prepared (perhaps by producing better vaccines or by developing other measures). The “bullet points” on your page 3 are really some of the arguments proposed by scientists like Lisa Jackson, Tom Jefferson, and some others for why the vaccines might not be working. But the key point in the article is the rather technical point that the best tests of effectiveness of the vaccines (randomized, double-blind, placebo-controlled) have never actually been carried out.

3. This is a good point. There is also undoubtedly a certain appeal connected to stories about courageous pioneers in science “bucking the received wisdom” and staking out controversial new positions. I think that angle of the story is probably one reason why this article appeared in a general-audience periodical like the Atlantic.

4. What you are saying here is only partly true. Maybe the most perplexing aspect of this whole story is why medical professionals are opposed to placebo-controlled trials for flu vaccines “for ethical reasons.” You do not say enough about this, but it is a key point in this whole discussion. The first ethical principle of medicine (from the Hippocratic oath) is “do no harm.” If, as a doctor, you give someone a placebo in a trial of a vaccine that turns out to be effective, haven’t you done harm if they get sick and possibly die from the illness? To justify that action, you have to argue that the greater good for society of doing the vaccine trial outweighs the concerns of individual patients. But that is not the way doctors are trained to think and I am not sure I would want their focus on the individual to change(!) This seems to be a real ethical dilemma to me!

Content:

Form: